OpenStudy (anonymous):

I'm finishing up a problem and don't understand how the answer is what it is... Here goes... -28/3Y = -28 (the step says to divide both sides by -28/3) and the answer should be Y=3 HOW?????

zepdrix (zepdrix):

Is this what it looks like?\[\large \frac{-28}{3}y=-28\]The \(\large y\) is not in the denominator of that fraction, correct?

OpenStudy (anonymous):

sorry, was answering someones post too... and yes correct it is how you put it up there...

zepdrix (zepdrix):

If we had\[\large 4y=4\] We would divide by 4 to isolate the \(\large y\) term, yes? Basically what we did was, we divided by the size of the `coefficient` on \(\large y\). So it's clear that the -28/3's will cancel on the y side right? We need to revisit what happens when we divide by a fraction. (That will tell us what's happneing on the right side of the equation).

zepdrix (zepdrix):

\[\large \dfrac{\cancel{\dfrac{-28}{3}}y}{\left(\cancel{\dfrac{-28}{3}}\right)}=\dfrac{-28}{\left(\dfrac{-28}{3}\right)}\] Understand that these will cancel on the left since they're the same value?

OpenStudy (anonymous):

yes, exactly....

OpenStudy (anonymous):

oh my...... and when I put it in the calculator the way you have it written... it comes out 3

zepdrix (zepdrix):

When we divide by a fraction, we can `flip it` and rewrite the operation as multiplication. Here's an example.\[\large \frac{\left(\dfrac{1}{2}\right)}{\left(\dfrac{3}{4}\right)} \qquad = \qquad \frac{1}{2}\cdot\frac{4}{3}\] See how I `flipped`, (took the reciprocal) of the bottom one?

zepdrix (zepdrix):

Oh yes, if you feel more comfortable just using a calculator ~ Then make sure you use some brackets.

OpenStudy (anonymous):

for some reason, every way I put it in before I would get the answer 1/3...... and that is wrong!!!!

OpenStudy (anonymous):

yep, I used the brackets, but I misplaced them!!!! Thanks for your help... medal is on the way if it will let me give it to you!!!! (now I can move on to the resat of my work)!THANKS

zepdrix (zepdrix):

\[\large \frac{-28}{\left(\dfrac{-28}{3}\right)} \qquad = \qquad -28\cdot \frac{3}{-28}\] See how I flipped the bottom fraction and rewrote it as multiplication using that rule?

zepdrix (zepdrix):

heh np XD